OpenStudy (anonymous):

Help Please What is the sum of the angle measures of a 36-gon? a.)6,480 b.)6,300 c.)6,840 d.)6,120

OpenStudy (solomonzelman):

for any shape with n sides, \(\large\color{blue}{ (n-2)(180) }\)

OpenStudy (solomonzelman):

Triangle = \(\large\color{blue}{ (3-2)(180)=180 }\) Square (or any 4-gon) = \(\large\color{blue}{ (4-2)(180)=360 }\) .... 36-gon =\(\large\color{blue}{ (36-2)(180)=? }\) you tell me

OpenStudy (anonymous):

6,120??

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

thank you so much :)